I guess I am just confused with mostly how to get the Ending Inventory column. Any help is appreciatedEquivalent Units Calculations-FIFO Method Ferris Corporation makes a powdered rug shampoo in two sequential departments, Compounding and Drying. Materials are added at the beginning of the process in the Compounding Department. Conversion costs are added evenly throughout each process. Ferris uses the FIFO method of process costing. In the Compounding Department, beginning work in process was 4,000 pounds (70% processed), 37,000 pounds were started in process, 36,000 pounds transferred out, and ending work in process was 60% processed Calculate equivalent units for the Compounding Department for March 2016 Ferris Corporation Flow of Units and Equivalent Units Calculation, March 2016 Equivalent Units %in March Direct Materials %in March Conversion Costs Complete/Transferred 4,000 1,200 Beginning inventory Started & completed 32,000 100 32,000 100 32,000 Ending 4,000 100 60 inventory Total
Answer #1

Computation of ending inventory in FIFo method in march

__________________________________________________________

Given:

Begining work in process = 4000 pounds

Started in process = 37000 pounds

Transferred out = 36000 pounds

Ending inventory = 60%

Ending inventory = begining work in process units

+ Started in process units

- transferred out units

= 4000+37000-36000

= 5000 pounds

Ending inventory = 5000 pounds

Total equipment units =begning work in process +

Started and completed +

Closing inventory

= 4000 +32000+ 5000

= 41000 pounds

Equivalent units for direct material

___________________________________

Closing inventory for

direct material = it is given that direct material

Added at the beginning of the

Process so take it as 100%

For equivalent units

Ending inventory = 5000 pounds for direct material

Total equivalent units

For direct material = 32000+5000

= 37000 pounds

Equivalent units for convertion cost

____________________________________

Ending inventory = 5000pounds *60%

= 3000 pounds

Total equivalent units

For convertion cost = 1200+32000+3000

= 36200 pounds